Use incremental analysis to select the best alternatives using Benefit – Cost ration

16. Use incremental analysis to select the best alternatives using Benefit - Cost ration Cost 4000 2000 7330 47000 8730 1000 1340 9000 9000 10000 9500 6000 Pw(benefits) Useful life is 20years Interest 6%
Answer #1

BCR= PV Benefits/ Pv Cost

Where PV Benefits = Present value of benefits

Pv Cost = Present value of cost

A . Pw of benefits = 7330 (p/f, 6%,20 )=7330 * .312 = 2285.53

BCR = PV Benefits/ PV Cost

= 2285.53 / 4000

= 0.5713

B . Pw of benefits = 47000 ( P/F , 6 % ,20 ) = 47000 * 0.312 = 14664

BCR = PV Benefits/ PV Cost

= 14664 / 2000

= 7.332 > 1

C. Pw of benefits = 8730 ( P/F , 6 % ,20 ) = 8730 * 0.312 = 2723.76

  BCR = PV Benefits/ PV Cost

= 2723.76 / 6000

= 0.45396

D. Pw of benefits = 1340 ( P/F , 6 % ,20 ) = 1340 * 0.312 = 418.08

BCR = PV Benefits/ PV Cost

= 418.08 / 1000

= 0.41808

E. Pw of benefits = 9000 ( P/F , 6 % ,20 ) = 9000 * 0.312 =2808

  BCR = PV Benefits/ PV Cost

= 2808 / 9000

= 0.312

F. Pw of benefits = 9500 ( P/F , 6 % ,20 ) = 9500 * 0.312 =2964

BCR = PV Benefits/ PV Cost

= 2964 / 10000

= 0.2964

Since, B / C of the incremental alternative is greater than one,

choose the higher first cost alternative which is alternative B